the sum of the ages of a father and his son is 41years If the product of their ages is 210 find their ages
Answer by Alan3354(69443)

Try pairs of factors of 210.
Find a pair with a sum of 41.
-------------
1*210 NG
2*105 NG
etc
============
The other tutor's solution:
et x be the father's age and y be the son's age, then
:
1) x + y = 41
:
2) xy = 210
:
solve equation 1 for x and substitute for x in equation 2
:
x = 41 - y
:
(41-y)y = 210
:
41y -y^2 = 210
:
y^2 -41y +210 = 0
:
(y-35) * (y-6) = 0 ****** At this point, he found a pair of factors of 210 with a sum of 41.